a dry cell when short-circuited will furnish about 30 amperes for a brief time. If its emf is 1.5 volts, ehat is its internal resistance? an ordinary household electric lamp takes about 1 amp will it be safe to connect it directly to the dry cell? Why?
1
Expert's answer
2013-08-06T08:27:29-0400
I=E/(r+R), where
I current
E EMF, r is internal resistance, R is resistance of whole circuit.
If the circuit is short-circuited then R will be 0 ohm
r=E/i
r=1.5/30=0.05 ohm
P=I*E
P(if R=0)=1.5*30=45 w(on internal resistance)
P(if R not 0, light bulb resistance, for example)=1*1.5=1.5 w - is much less
than 45 w
but P(r)=1*0.05=0.05 w(power on internal resistance, with connected light bulb)